If a ball is tossed up vertically from the ground at 20m/s, what is it’s velocity 1 second before reaching maximum height?
Answer #1

Gravitational acceleration = g = -9.81 m/s2

Initial velocity of the ball = V0 = 20 m/s

Velocity of the ball at maximum height = V1 = 0 m/s

Time taken to reach maximum height = t1

V1 = V0 + gt1

0 = 20 + (-9.81)t1

t1 = 2.038 sec

Velocity of the ball 1 second before reaching maximum height = V2

Time the ball has been in air = t2

t2 = t1 - 1

t2 = 2.038 - 1

t2 = 1.038 sec

V2 = V0 + gt2

V2 = 20 + (-9.81)(1.038)

V2 = 9.81 m/s

Velocity of the ball 1 second before reaching maximum height = 9.81 m/s